OpenStudy (anonymous):

How do I find a cartesian equation for the curve r=3/(1+sin(theta)) ?

OpenStudy (shamim):

\[r=\frac{ 3 }{ 1+\sin \theta }\]

OpenStudy (shamim):

\[or, r+r \sin \theta=3\]

OpenStudy (shamim):

i think u know \[x=r \cos \theta\]

OpenStudy (shamim):

\[y= r \sin \theta\]

OpenStudy (shamim):

\[x ^{2}+y ^{2}=r ^{2}\]

OpenStudy (shamim):

so our equation will b\[r+y=3\]

OpenStudy (shamim):

i think u r now able to simplify it

OpenStudy (anonymous):

so from here, would i move it to r=3-y then square both sides to r^2=9-6y+y^2 ?

OpenStudy (shamim):

ya

OpenStudy (shamim):

put r^2=x^2+y^2

OpenStudy (anonymous):

oh! ok, i get it now! Thank you!

OpenStudy (shamim):

welcome